A pendulum of length 28cm oscillates such that its string makes an angle of 300 from the vertical when it is at one of the extreme positions. Find the ratio of the distance to displacement of the bob of the pendulum when it moves from one extreme position to other.


Grade:9

2 Answers

Ritender Singh AskiitiansExpert-IIT
13 years ago

First try to find out displacement:

as the angle between from one extreme position to other extreme position is 60 degree

So, displacement = 28 cm (Calculated by geometry)

Now on distance:

 = (pie/6)*28 cm

So, Ratio = (Distance/Displacement) = (pie/6)

= 0.523 Ans

Prakash Kumar
7 years ago
Displacement= 2*r*sintheta
Distance= pi* r*2theta/180.     Hence the ratio will be = pi* r*2theta/180 / 2*r*sintheta =22:21